OpenStudy (anonymous):

will give 2 medals Which expression is equivalent to one over fivem − 20? one over five(m − 4) one over five(m − 100) 5(m − 4) 5(m − 100)

OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 4 points) Write an expression that is equivalent to three over five (3y + 15). three over fivey + 9 three over five y + 15 nine over fivey + 9 nine over fivey + 15 Question 3 (Multiple Choice Worth 4 points) Write an expression for the perimeter of the triangle shown below: A triangle is shown with side lengths labeled 0.3x plus 24, 5x, and 0.2x minus 15. 10x − 9 10x − 39 5.5x − 9 5.5x + 9 Question 4 (Multiple Choice Worth 4 points) Monica purchased adult and youth tickets for the fair. She bought x adult tickets for $28 each and y youth tickets for $15. Write an expression that can be used to show the total cost, C, of all tickets. C = (28 + 15)x C = (28 + 15)(x + y) C = 28x + 15y C = 28y + 15x Question 5 (Multiple Choice Worth 4 points) Ella purchased a game that was on sale for 12% off. The sales tax in her county is 6%. Let y represent the original price of the game. Write an expression that can be used to determine the final cost of the game. y − 0.12y 0.06(0.88y) 1.06(0.88y) y − 0.88y + 0.06y

OpenStudy (anonymous):

please help me @DanJS your my only hope

OpenStudy (danjs):

These are kinda awkward. Judging by the answers for the first one, this is the expression in the problem. \[\frac{ 1 }{ 5m-20 }\] Factor a 5 out in the denominator

OpenStudy (anonymous):

but do you know what its equivalent to i got 150

OpenStudy (anonymous):

not for what its equivalent to but in the options

OpenStudy (danjs):

It doesn't have a single value. There is a variable in the expression. All you can do is alter it or simplify. Are the problems actually written out in english like this? It is difficult to tell what is numerator , denominator, or quantities. I can tell because of the multiple choices, but it is not clear from the problem.

OpenStudy (danjs):

\[\frac{ 1 }{ 5m-20 } = \frac{ 1 }{ 5(m-4) }\]

OpenStudy (danjs):

For the second... \[\frac{ 3(3y+15) }{ 5 }\] The(3y+5) i am assuming is in the numerator, by what the answers are. factor a 3 out of the parenthesis in this one.

OpenStudy (anonymous):

which one is this for

OpenStudy (danjs):

The perimeter of the triangle, is the sum of the 3 given sides. Add up like terms.

OpenStudy (anonymous):

what do i have to do know